(nt, K) and two baryons (p, n). 1) [5 pts] Sketch a Quark compositions of two mesons studying for my final, help
Answer #1

There are three quarks: up, down and strange and each has a corresponding antiquark of opposite charge.

Charge/e = +2/3 Charge/e = -1/3
up u down d
charm c strange s
top t bottom b
Baryon quark composition charge/e
neutron udd +2/3 -1/3 - 1/3 = +1
proton uud +2/3 +2/3 + 1/3 = +1
Meson quark composition charge/e
pion \pi ^{+} u\bar{d} +2/3 + 1/3 = +1
kaon K^{-} s\bar{u} -2/3 - 1/3 = -1

If any doubt feel free to comment